OpenStudy (anonymous):

Which grouping of numbers gives the expression a value of 15? i chose 6+3 x (18/9)-3 and got it wrong???

OpenStudy (anonymous):

that actually equals 9 my friend

OpenStudy (anonymous):

6+3= 9, times 2 and then -3

OpenStudy (anonymous):

you're supposd to use PEMDAS

OpenStudy (mathstudent55):

Do you have choices?

OpenStudy (anonymous):

perenthesis first

OpenStudy (anonymous):

18/9 = 2

OpenStudy (anonymous):

Yes, one sec

OpenStudy (anonymous):

now you have 6+3 * 2 -3

OpenStudy (anonymous):

next you multiply now you have 6+ 6 -3

OpenStudy (anonymous):

then add now you have 12 -3

OpenStudy (anonymous):

now you have 9

OpenStudy (anonymous):

6+(3x18/9)-3, 6+(3x18)/9-3, 6+3x18/(9-3)

OpenStudy (mathstudent55):

\(6+3 \times \color{red}{(18 \div 9)}-3\) Parentheses first: \(= 6+3 \times \color{red}{2}-3\) \(= 6+\color{red}{3 \times 2}-3\) Then the multiplication: \(= 6+\color{red}{6}-3\) Then additions and subtractions in the order they appear from left to right: \(= 6+6-3\) \(= 12 - 3\) \(= 9\)

OpenStudy (mathstudent55):

PEMDAS is the order of operations: P - Parentheses E - Exponents MD - Multiplications and divisions in the order they appear from left to right AS - Additions and subtractions in the order they appear from left to right

OpenStudy (anonymous):

6+3x18/(9-3) 6+3x18/(6) 6+54 / 6 6+9 15

OpenStudy (anonymous):

15

OpenStudy (mathstudent55):

@GBEiswhatmytitlesay Try this one yourself. What do you get? 6+(3x18)/9-3

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

18 x 3 is 54...sooo 6+54/9-3 54/9= 6 & 6+6-3...6+6 is 12 and -3 is 9

OpenStudy (mathstudent55):

Here it is step-by-step. It's exactly what you did.